June 17, 2015 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, June 17, 2015
11:00 am, EDT
JLab: CEBAF Center, F326

Communication Information

Remote Connection

You can connect using BlueJeans using the meeting number 630 804 895 .       (Click "Expand" to the right for more details -->):

  1. Make sure you have created a BlueJeans account via your JLab CUE account using this link:

  2. Meeting ID: 630804895
    • (you may need to type this in, depending how you connect)

  3. If connecting via Web Browser: click this link (no passcode is needed):

  4. If connecting via iOS or Android App:
    • Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting

  5. If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"

  6. If connecting via Polycom unit:
    • Dial or bjn.vc
    • Enter meeting ID above
    • Use *4 to unmute


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2015 on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2015/ .


  1. Announcements
  2. Subdetector Reports
  3. Simulations
  4. Calibration Processing
  5. AOB


Attending: Sean (NU); Simon, Beni, Luke, Adesh, Eric, Lubomir, Greg (JLab); Naomi, Will M. (CMU)

  1. CDC - Simon reported that has some new alignment constants from the cosmic runs in the spring, which don't have some of the unusual features of his previous constants based on data taken last fall. He is checking them out in preparation for putting them in the CCDB, and will compare these results to Mike Staib's. He also updated the CDC energy scale constants to fix some run-dependence of the dE/dx in the spring data and to bring it in line with what is expected from MC. Naomi looked at this data and thought it could be due to the threshold values being changed while appropriate values were being determined.
  2. FDC - Lubomir is working on implementing a new method of calculating the centroid on the cathode strips, to improve resolution and fix the systematics in FDC hit position that he has been studying. The status of FDC gains was briefly discussed, and he will ask Vlad to take a look at the per-channel gain calibrations with a run from the spring.
  3. FCAL - Adesh reported that he has made a lot of progress in studying the FCAL calibrations after fixing a problem with his code. He sees a difference in the pi0 width depending on the different data modes and running conditions. He is compiling these results to present at the next calorimetry meeting. [n.b. tomorrow's meeting is canceled]. Elton has also taken a look at using electrons for gain calibration, and Adesh is comparing these with his results.
  4. BCAL - Will reported that George is working on calibrating the effective velocities and time offsets, and Noemi is calibrating the TDC timewalks. Neither has results to report at the moment.
  5. Start Counter - Eric reported that they have concluded that there is not enough beam data to perform an accurate calibration of the propogation and attenuation corrections for each paddle, so for now they are re-analyzing their bench data, and preparing new constants and the reconstruction code that will use these constants. The constants obtained from the be bench and beam data look generally consistent. They are also refining and automating their procedures. Sean stated that it would be good to see updated results in the next calibration meeting, since several people are looking at reconstruction issues in the forward region, where improved start counter times could be useful.
  6. TOF - Beni reported that for his calibrations he used several runs, but found that one of them had TOF times shifted by 4 ns. He thinks the run is 3079, but will check. Once the run is found, Sean or Simon can updated the CCDB to fix this shift.
  7. Simulations - It was decided last week to wait for a report from Tegan about his updated BCAL simulation code before generating more simulations. This report is expected next week, after the IU workshop is concluded.
  8. Calibration Processing - Sean reported on the common calibration jobs he plans to run. He wrote a framework to run these jobs on the batch farm or interactively/via cronjob [intended for use in the online]. For the batch farm, he is using the SWIF workflow management tool to do most of the job management. He and Kei are now Users #1 and 2 of this tool, so much of the past week and a half was spent debugging problems with SWIF. As of yesterday, he has some simple jobs working, and is talking to Will McGinley and Mike Staib about integrating their calibrations into this workflow. He welcomed any other calibration plugins that people would like run, any runs that should be added to the list, and any other comments.